Zion product corporations have the following capital, structure, which it considers optimal:

Bonds, 8% (par value $1000) $ 400000

Preferred stock 200000

Common stock 300000

Retained earnings 300000

Total 1200000

Additional information:

-dividends on common stock and preferred stocks are currently $ 3 and $ 5 per share respectively and are expected to grow at constant rate of 5% for common stocks.

-market price of common stock is $ 30 and the preferred stock is selling at $ 40

-flotation cost on new issues of common stock and preferred stock are $ 2 and 4 respectively.

-the bond is sold out at par value.

-the interest on bonds is paid annually for 5 years and the company’s tax rate is 40%

Calculate the cost of new common stock and the weighted average cost of capital


1
Expert's answer
2021-03-29T11:42:18-0400

the cost of new common stock:

"\frac{3}{30-2}+0.05= 0.1571=15.71%"

the cost of new preferred stock:

"\frac{5}{40-4}+0.05= 0.1889=18.89%"

Market price of debt:

interest expense: "400,000*0.08=32,000"

Cost of debt: "0.08*0.6=0.048"

Market price: "32,000*\frac{(1-\frac{1}{(1+0.048)^5})}{0.048}+\frac{400,000}{(1+0.048)^5}=455,725"

Total capital: "300,000+200,000+300,000+455,725=1,255,725"

"WACC=\frac{300,000+300,000}{1,255,725}*15.71+\frac{200,000}{1,255,725}*18.89+\frac{455,725}{1,255,725}*0.048=10.53\%"